a pendulumof length 1.5 m and mass 3 kg is created

a. find time it takes the pendulum to first be virtical assuming release angle of 5 degrees from vertical and near suface of earth

b. find same as above, but on the moon

c. find time with release angle of 90 degrees near earths surface

d. find time it takes to first be virtical  assuming release angle of 5 degrees, near earth, but double the mass
Answer #1

g = 9.8 m/s^2   , L =1.5 m   , m = 3 kg


a) T = time period = 2*pi*[L/g]^0.5 = 2.45 s

     t = time reqiured = T/4 =0.61 s


b) on moon , g =9.8/9 = 1.088

         T' = 2*pi*[L/1.088]^0.5 = 7.35 s

        time   = T'/4 = 1.83 s



c) time = 0.61 s

d) t = 0.61